OpenStudy (anonymous):

Let F(x)= 4 + 5x and g(x) = 2x-1. Find f(g(x)) and g(f(x)) f(g(x))

OpenStudy (astrophysics):

f(g(x)) means you plug in the function g(x) wherever there is an x in f(x) so we have \[f(g(x)) = 4+(2x-1)\] now evaluate

OpenStudy (astrophysics):

\[f(g(x)) = 4+5(2x-1)\]

OpenStudy (anonymous):

f(2x-1)=4+10x-5

OpenStudy (astrophysics):

Similarly with g(f(x)) you plug in f(x) in function g(x)

OpenStudy (anonymous):

\[g(f(x))=2(4+5x)-1\]

OpenStudy (anonymous):

@Astrophysics right?

OpenStudy (astrophysics):

Yes but we should have left @nabiladh do that on their own to see if they understood the problem :)

OpenStudy (anonymous):

yes right,sorry if i done wrong i would delete it

OpenStudy (anonymous):

f(2x-1)=10x-1

OpenStudy (astrophysics):

Nope, what is 2*4?

OpenStudy (anonymous):

my bad (8+10x)-1

OpenStudy (astrophysics):

Yeah now keep simplifying

OpenStudy (anonymous):

-8-10x?

OpenStudy (astrophysics):

2(4+5x)-1 = 8+10x-1 = 10x+7

OpenStudy (anonymous):

so f(g(x)) equals 10x+7?

OpenStudy (astrophysics):

No...that's g(f(x))

OpenStudy (astrophysics):

Ok I think there's a lot of confusion because both were done at the same time so f(g(x)) is \[f(g(x)) = 4+5(2x-1) = 4+10x-5 = 10x-1\]

OpenStudy (astrophysics):

Then g(f(x)) is \[g(f(x)) = 2(4+5x)-1 = 8+10x-1 = 10x+7\]

OpenStudy (anonymous):

@Astrophysics i see now thanks confusing but i got it

OpenStudy (astrophysics):

Ok cool :)

OpenStudy (jmark):

f(g(x) = 4 + 5(2x-1) = 4 + 10x - 5 = 10x - 1. g(f(x))= 2(4+5x) - 1 = 8+10x-1 = 10x+7